# IDALC: A Semi-Supervised Framework for Intent Detection and Active Learning based Correction

### v1

@@ 2.1 Intent Class Detection
In dynamic and evolutionary real-world applications, various flexible adaptive intent class detection models are developed. [liao2023novel, kuzborskij2013n, scheirer2012toward, degirmenci2022efficient] work on incremental learning in a dynamic environment for evolving new classes in streams. sun2016online aims at the class evolution phenomenon of the emergence and disappearance of classes. Some advanced approaches are proposed to explore novel class detection problems in data streams. ECSMiner [masud2010classification] resolve streaming problems but the time delay assumption and fixed-size chunks are unrealistic in real scenarios. SAND [haque2016sand] is a semi-supervised framework to identify concept drifts. mullick2022framework identifies novel intents but does not apply any approach to retain the accuracy of known intents and does not handle system-rejected utterances. In different specialized contexts, researchers also work on intent phrase identification task [mullick2022fine, mullick2023novel, mullick2023exploring, mullick2022evaluation, mullick2024pointer, mullick2023intent, mullick2024intent, mullick2025text], various entity-extraction [mullick2024matscire, guha2021matscie, mullick2022using, mullick2021rte], opinion context detection [mullick2017generic, mullick2016graphical, mullick2018identifying, mullick2018harnessing, mullick2019d, mullick2017extracting] and others [mullick2024avenues, mullick2025introducing]. Based on the Frequent-Directions technique [liberty2013simple], SENC-MaS [mu2017streaming] is a matrix sketch method to detect new classes and update the classifier. SENCForest [mu2017classification] utilizes random trees as a single common core to identify new classes. But both these approaches rely on a large labelled initial training set and do not use unlabelled data. [cui2023novel, abdallah2016anynovel, zhou2022detecting] explore outlier detection approaches but most of them are applicable on images. [na2018dilof, zhan2021out, larson2019evaluation, yan2020unknown, zhou2022knn, firdaus2023multitask] detect new class in the form of outlier detection. However, these algorithms fail to identify new classes from rejected utterances and classify all. SEEN [zhu2020semi] uses both labeled and unlabelled data to detect emerging new classes and classify known classes. Zero-shot-OOD [tan2019out] proposes a new method to tackle the OOD detection task in low-resource settings, without adversely affecting performance on the few-shot ID classification. Aligner [zhu2024aligner2], CEA-Net [wu2024cea] explore joint multiple intent detection with slot filling task. park2024joint apply the joint-BERT model for domain-specific intent classification. SENNE [cai2019nearest] and IFSTC [xia2021incremental] handle streaming class classification problems. TARS [halder-etal-2020-task] identifies in-domain data and classifies them. We use SEEN, SENC-MaS, Zero-shot-OOD, SENNE, IFSTC, and TARS as baselines.
@@ 2.2 Active Learning based Correction
Active learning (AL) is a machine learning paradigm that aims to reduce the annotation effort by selectively choosing the most informative data points to label. Instead of randomly labeling large volumes of data, AL focuses on identifying uncertain or representative instances that are likely to improve the model’s performance when labeled. This is particularly useful in tasks like intent detection, where obtaining labeled data can be expensive and time-consuming, especially when dealing with large or diverse datasets. wang2020active proposes an active learning strategy, based on label error statistical methods to find critical instances. salama2022novel uses active learning for heterogeneous data. senay2020viraal, zhang2020deep apply active learning (AL) method for intent classification. yin2023clustering is a cluster-based active learning method works on data streams. But none of these works focused on accuracy improvements of intent classes using AL approaches. gao2020consistency presents a consistency-based semi-supervised AL framework and a simple pool-based AL selection metric to select data for labeling. But these models do not handle system-rejected samples and are useful for images, not in textual contexts. modAL2018 proposes various AL strategies to improve system accuracy. tian2011active develops a maximal marginal uncertainty-based AL model to maximize accuracy with minimal interactions. We use the last two approaches as baselines (MaxMU-AL and MinU-AL).
@@ 3 Dataset
We perform our experiments on three NLU benchmarked datasets - SNIPS [couckesnips:2018], ATIS [tur2010left] and Facebook Multi-lingual data [schuster2018cross].
@@ 4 Approach
Neural Model ($\mathcal{M}$): We use the JointBERT [chen2019bert] model as our learning model. The model performs intent classification. We do not use the auxiliary task of slot-filling. We use English uncased BERT-Base [^1] We explore different models but it performs the best model [devlin2019bert] as the base model for JointBERT. In the case of other languages, we use the bert-base-multilingual-uncased model. For the neural model, we perform a hyperparameter search and report the results of the settings that achieve the best results and then fix the same for all the models. The batch size is 16, the number of epochs is 10, the optimization algorithm used is AdamW and the learning rate is 5e55e-5 with cross-entropy as the loss function.
@@ 5.1 Intent Detection
Out-of-Domain Sample Detection Algorithms (OOD-SDA): We consider three algorithms for detecting out-of-domain samples. i) Softmax Prediction Probability (MSP) [hendrycks2018baseline], ii) Deep Open Classification (DOC) [shu2017doc] method and iii) Local Outlier Factor (LOF) [10.1145/342009.335388]. For MSP, we use the minimum prediction score on the training set to decide the threshold and report for the best threshold. We experiment on all English datasets as shown in Table 2. DOC performs the best in terms of both accuracy (Acc) and macro-f1(F1), so we use the DOC model in our pipeline and for different experiments.
@@ 5.1.1 Competing Baselines
SEEN [zhu2020semi] is designed specifically for streaming intent detection, where new labels can appear continuously in the incoming data. It combines two key components: (i) SEEN-Forest, an ensemble of random trees that serves as an anomaly detector, and (ii) SEEN-LP, a graph-based label propagation module for semi-supervised classification of queries. This hybrid design makes SEEN particularly relevant for dialog systems, as it explicitly addresses the challenge of detecting novel intents with minimal human supervision. In our setup, we initialize their model with a small amount of labeled data in the initial stream (stream 0), incrementally update it with unlabeled data from the next stream (stream 1), and finally evaluate it on the held-out test set. The complexity of label propagation is $\mathcal{O}(n_{\text{train}}^{2})$ in dense graphs or $\mathcal{O}(n_{\text{train}})$ when sparsified, while training decision trees costs $\mathcal{O}(n_{\text{train}}\log n_{\text{train}})$.
@@
SENC-MaS [mu2017streaming] builds compact matrix sketches that summarize the distribution of known-class data and then leverages these sketches for both classification and novelty detection. By operating on compressed sketches rather than the raw data, it prioritizes efficiency, making it well-suited for high-throughput dialog streams where retraining large neural models on every update is infeasible. In our experiments, the algorithm incrementally updates its sketches with streaming unlabeled data, after which new queries are either classified or flagged as novel. The update complexity is $\mathcal{O}(n_{\text{train}}\,d)$, reflecting its linear dependence on training data and hidden dimension.
@@
SENNE [cai2019nearest] is a nearest-neighbor ensemble approach tailored for emerging-class detection, especially in scenarios where new intents are closely related to existing ones. This characteristic is important for dialog systems, as subtle distinctions (e.g., PlayMusic vs. AddToPlaylist) can critically impact performance. SENNE works by constructing ensembles of nearest-neighbor classifiers for each known class and checking whether a query falls within or outside the decision boundary of these ensembles. This makes it particularly strong in handling borderline cases where unseen intents resemble existing categories. The computational cost is $\mathcal{O}(n_{\text{train}}\,d)$ at inference.
@@
IFSTC [xia2021incremental] frames intent detection as a textual entailment problem, which enables few-shot classification of new intents without retraining on the entire dataset. This is especially relevant for IDALC, as both approaches aim to generalize quickly from small amounts of annotated data. IFSTC employs a hybrid strategy: it initializes models with XNLI pretraining and then fine-tunes them using only a few labeled examples of new intents. This combination of cross-lingual pretraining and few-shot adaptation makes it robust across languages and efficient in data-scarce settings. Its fine-tuning cost is $\mathcal{O}(e\,n_{\text{train}}\,L^{2}\,d\,\ell_{\text{BERT}})$.
@@
Zero-Shot-OOD [tan2019out] is built on an OOD-resistant prototypical network, which learns compact clusters of known-class examples and rejects queries that fall outside them. The method directly benchmarks IDALC’s OOD detection capabilities, since both target the challenge of handling unseen intents with limited supervision. In our experiments, prototypes were constructed using meta-learning episodes, and incoming queries were tested for membership within known clusters. Queries sufficiently distant from all prototypes were flagged as out-of-distribution. Training scales as $\mathcal{O}(e\,n_{\text{train}}\,d)$ and inference as $\mathcal{O}(k\,d)$ per query.
TARS [halder-etal-2020-task] reformulates text classification as a label-text matching task, where each input query is paired with an intent label and the model decides whether they match. This approach is highly relevant for our evaluation, as it probes whether intent names alone can serve as sufficient supervision. Unlike IDALC, which uses a more elaborate loop of OOD detection, clustering, and annotation, TARS relies directly on semantic alignment between queries and labels. For our benchmarks, we trained TARS models on dialog datasets using BERT-base for English and multilingual BERT for non-English data. The training cost is $\mathcal{O}(e\,n_{\text{train}}\,L^{2}\,d\,\ell_{\text{BERT}})$, with inference requiring a single forward pass per label.
@@ 5.2.1 Baselines
We apply Maximum Marginal Uncertainty (MaxMU-AL) with minimum interaction [tian2011active] and Minimum Uncertainty (MinU-AL) based Active Learning model [modAL2018] on top of intention detection (ID(1)+KM) output. We use ID(1)+KM, Joint Bert Neural Model ($\mathcal{M}$), MaxMU-AL, and MinU-AL as baselines.
@@ 5.5 Real World Scenarios
(i) In practical applications, user queries often contain multiple intents. To evaluate our approach in such cases, we used two benchmark datasets for multi-intent detection: Mix-SNIPS [qin2020agif] and Mix-ATIS [qin2020agif]. These datasets consist of queries with multiple intent labels. We extended the IDALC architecture to support sentence-level multi-label detection, enabling the system to predict multiple intents for a single query. The modified model achieved 76.45% accuracy on Mix-SNIPS and 70.37% on Mix-ATIS, showing its effectiveness and robustness in handling multi-intent queries.
@@ 8 Limitations
While our study shows promising results, it has several limitations that open important directions for future work: (1) ID-ALC relies on human-in-the-loop feedback, which improves accuracy but requires costly manual annotation; we plan to explore zero-shot adaptation using self-learning and large language models. (2) The framework currently supports only text, whereas real-world applications often involve voice, images, or mixed modalities; we will scale datasets and replace classifiers with multimodal models such as SDIF-DA [huang2023sdifda], MulT [tsai-etal-2019-multimodal], MISA [hazarika2020misa], and MAG-BERT [rahman-etal-2020-integrating]. (3) Our experiments are limited to English, Thai, and Spanish, leaving out code-mixed or code-switched settings; extending to such cases will require multilingual models with alignment across framework components. (4) We also plan to evaluate low-resource languages and domain-specific datasets (e.g., education, healthcare, law) to test adaptability beyond general-purpose intent detection. (5) While the current work focuses on optimizing performance with minimal annotation, future extensions will address the interpretability of IDALC, as understanding which features or interactions drive clustering and adaptation can improve trust, usability, and applicability in expert-facing domains.

### v2

@@ 2.1 Intent Class Detection
In dynamic and evolutionary real-world applications, various flexible adaptive intent class detection models are developed. [20, 18, 50, 8] work on incremental learning in a dynamic environment for evolving new classes in streams. [54] aims at the class evolution phenomenon of the emergence and disappearance of classes. Some advanced approaches are proposed to explore novel class detection problems in data streams. ECSMiner [22] resolve streaming problems but the time delay assumption and fixed-size chunks are unrealistic in real scenarios. SAND [14] is a semi-supervised framework to identify concept drifts. [41] identifies novel intents but does not apply any approach to retain the accuracy of known intents and does not handle system-rejected utterances. In different specialized contexts, researchers also work on intent phrase identification task [37, 43, 44, 38, 26, 36, 33, 42], various entity-extraction [28, 12, 39, 25], opinion context detection [35, 32, 29, 31, 40, 30] and others [34, 27]. Based on the Frequent-Directions technique [21], SENC-MaS [24] is a matrix sketch method to detect new classes and update the classifier. SENCForest [23] utilizes random trees as a single common core to identify new classes. But both these approaches rely on a large labelled initial training set and do not use unlabelled data. [6, 1, 66] explore outlier detection approaches but most of them are applicable on images. [45, 64, 19, 62, 67, 10] detect new class in the form of outlier detection. However, these algorithms fail to identify new classes from rejected utterances and classify all. SEEN [68] uses both labeled and unlabelled data to detect emerging new classes and classify known classes. Zero-shot-OOD [55] proposes a new method to tackle the OOD detection task in low-resource settings, without adversely affecting performance on the few-shot ID classification. Aligner [69], CEA-Net [60] explore joint multiple intent detection with slot filling task. [46] apply the joint-BERT model for domain-specific intent classification. SENNE [3] and IFSTC [61] handle streaming class classification problems. TARS [13] identifies in-domain data and classifies them. We use SEEN, SENC-MaS, Zero-shot-OOD, SENNE, IFSTC, and TARS as baselines.
@@ 2.2 Active Learning based Correction
Active learning (AL) is a machine learning paradigm that aims to reduce the annotation effort by selectively choosing the most informative data points to label. Instead of randomly labeling large volumes of data, AL focuses on identifying uncertain or representative instances that are likely to improve the model’s performance when labeled. This is particularly useful in tasks like intent detection, where obtaining labeled data can be expensive and time-consuming, especially when dealing with large or diverse datasets. [59] proposes an active learning strategy, based on label error statistical methods to find critical instances. [49] uses active learning for heterogeneous data. [52, 65] apply active learning (AL) method for intent classification. [63] is a cluster-based active learning method works on data streams. But none of these works focused on accuracy improvements of intent classes using AL approaches. [11] presents a consistency-based semi-supervised AL framework and a simple pool-based AL selection metric to select data for labeling. But these models do not handle system-rejected samples and are useful for images, not in textual contexts. [7] proposes various AL strategies to improve system accuracy. [56] develops a maximal marginal uncertainty-based AL model to maximize accuracy with minimal interactions. We use the last two approaches as baselines (MaxMU-AL and MinU-AL).
@@ 3 Dataset
We perform our experiments on three NLU benchmarked datasets - SNIPS [5], ATIS [58] and Facebook Multi-lingual data [51].
@@ 4 Approach
Neural Model ($\mathcal{M}$): We use the JointBERT [4] model as our learning model. The model performs intent classification. We do not use the auxiliary task of slot-filling. We use English uncased BERT-Base [^1] We explore different models but it performs the best model [9] as the base model for JointBERT. In the case of other languages, we use the bert-base-multilingual-uncased model. For the neural model, we perform a hyperparameter search and report the results of the settings that achieve the best results and then fix the same for all the models. The batch size is 16, the number of epochs is 10, the optimization algorithm used is AdamW and the learning rate is 5e55e-5 with cross-entropy as the loss function.
@@ 5.1 Intent Detection
Out-of-Domain Sample Detection Algorithms (OOD-SDA): We consider three algorithms for detecting out-of-domain samples. i) Softmax Prediction Probability (MSP) [16], ii) Deep Open Classification (DOC) [53] method and iii) Local Outlier Factor (LOF) [2]. For MSP, we use the minimum prediction score on the training set to decide the threshold and report for the best threshold. We experiment on all English datasets as shown in Table 2. DOC performs the best in terms of both accuracy (Acc) and macro-f1(F1), so we use the DOC model in our pipeline and for different experiments.
@@ 5.1.1 Competing Baselines
SEEN [68] is designed specifically for streaming intent detection, where new labels can appear continuously in the incoming data. It combines two key components: (i) SEEN-Forest, an ensemble of random trees that serves as an anomaly detector, and (ii) SEEN-LP, a graph-based label propagation module for semi-supervised classification of queries. This hybrid design makes SEEN particularly relevant for dialog systems, as it explicitly addresses the challenge of detecting novel intents with minimal human supervision. In our setup, we initialize their model with a small amount of labeled data in the initial stream (stream 0), incrementally update it with unlabeled data from the next stream (stream 1), and finally evaluate it on the held-out test set. The complexity of label propagation is $\mathcal{O}(n_{\text{train}}^{2})$ in dense graphs or $\mathcal{O}(n_{\text{train}})$ when sparsified, while training decision trees costs $\mathcal{O}(n_{\text{train}}\log n_{\text{train}})$.
@@
SENC-MaS [24] builds compact matrix sketches that summarize the distribution of known-class data and then leverages these sketches for both classification and novelty detection. By operating on compressed sketches rather than the raw data, it prioritizes efficiency, making it well-suited for high-throughput dialog streams where retraining large neural models on every update is infeasible. In our experiments, the algorithm incrementally updates its sketches with streaming unlabeled data, after which new queries are either classified or flagged as novel. The update complexity is $\mathcal{O}(n_{\text{train}}\,d)$, reflecting its linear dependence on training data and hidden dimension.
@@
SENNE [3] is a nearest-neighbor ensemble approach tailored for emerging-class detection, especially in scenarios where new intents are closely related to existing ones. This characteristic is important for dialog systems, as subtle distinctions (e.g., PlayMusic vs. AddToPlaylist) can critically impact performance. SENNE works by constructing ensembles of nearest-neighbor classifiers for each known class and checking whether a query falls within or outside the decision boundary of these ensembles. This makes it particularly strong in handling borderline cases where unseen intents resemble existing categories. The computational cost is $\mathcal{O}(n_{\text{train}}\,d)$ at inference.
@@
IFSTC [61] frames intent detection as a textual entailment problem, which enables few-shot classification of new intents without retraining on the entire dataset. This is especially relevant for IDALC, as both approaches aim to generalize quickly from small amounts of annotated data. IFSTC employs a hybrid strategy: it initializes models with XNLI pretraining and then fine-tunes them using only a few labeled examples of new intents. This combination of cross-lingual pretraining and few-shot adaptation makes it robust across languages and efficient in data-scarce settings. Its fine-tuning cost is $\mathcal{O}(e\,n_{\text{train}}\,L^{2}\,d\,\ell_{\text{BERT}})$.
@@
Zero-Shot-OOD [55] is built on an OOD-resistant prototypical network, which learns compact clusters of known-class examples and rejects queries that fall outside them. The method directly benchmarks IDALC’s OOD detection capabilities, since both target the challenge of handling unseen intents with limited supervision. In our experiments, prototypes were constructed using meta-learning episodes, and incoming queries were tested for membership within known clusters. Queries sufficiently distant from all prototypes were flagged as out-of-distribution. Training scales as $\mathcal{O}(e\,n_{\text{train}}\,d)$ and inference as $\mathcal{O}(k\,d)$ per query.
TARS [13] reformulates text classification as a label-text matching task, where each input query is paired with an intent label and the model decides whether they match. This approach is highly relevant for our evaluation, as it probes whether intent names alone can serve as sufficient supervision. Unlike IDALC, which uses a more elaborate loop of OOD detection, clustering, and annotation, TARS relies directly on semantic alignment between queries and labels. For our benchmarks, we trained TARS models on dialog datasets using BERT-base for English and multilingual BERT for non-English data. The training cost is $\mathcal{O}(e\,n_{\text{train}}\,L^{2}\,d\,\ell_{\text{BERT}})$, with inference requiring a single forward pass per label.
@@ 5.2.1 Baselines
We apply Maximum Marginal Uncertainty (MaxMU-AL) with minimum interaction [56] and Minimum Uncertainty (MinU-AL) based Active Learning model [7] on top of intention detection (ID(1)+KM) output. We use ID(1)+KM, Joint Bert Neural Model ($\mathcal{M}$), MaxMU-AL, and MinU-AL as baselines.
@@ 5.5 Real World Scenarios
(i) In practical applications, user queries often contain multiple intents. To evaluate our approach in such cases, we used two benchmark datasets for multi-intent detection: Mix-SNIPS [47] and Mix-ATIS [47]. These datasets consist of queries with multiple intent labels. We extended the IDALC architecture to support sentence-level multi-label detection, enabling the system to predict multiple intents for a single query. The modified model achieved 76.45% accuracy on Mix-SNIPS and 70.37% on Mix-ATIS, showing its effectiveness and robustness in handling multi-intent queries.
@@ 8 Limitations
While our study shows promising results, it has several limitations that open important directions for future work: (1) ID-ALC relies on human-in-the-loop feedback, which improves accuracy but requires costly manual annotation; we plan to explore zero-shot adaptation using self-learning and large language models. (2) The framework currently supports only text, whereas real-world applications often involve voice, images, or mixed modalities; we will scale datasets and replace classifiers with multimodal models such as SDIF-DA [17], MulT [57], MISA [15], and MAG-BERT [48]. (3) Our experiments are limited to English, Thai, and Spanish, leaving out code-mixed or code-switched settings; extending to such cases will require multilingual models with alignment across framework components. (4) We also plan to evaluate low-resource languages and domain-specific datasets (e.g., education, healthcare, law) to test adaptability beyond general-purpose intent detection. (5) While the current work focuses on optimizing performance with minimal annotation, future extensions will address the interpretability of IDALC, as understanding which features or interactions drive clustering and adaptation can improve trust, usability, and applicability in expert-facing domains.